OpenStudy (anonymous):

what is the derivative of ln(ln x^2)

OpenStudy (tkhunny):

You've a chain rule in your future. Let's see your first attempt.

OpenStudy (anonymous):

can i use u-substitution?

OpenStudy (anonymous):

ln(u)= du/u?

OpenStudy (anonymous):

yes, that will work.

OpenStudy (anonymous):

so ln(ln x^2)=(d/dx (ln x^2)) / ln(x^2)

OpenStudy (anonymous):

2x/x^2/ ln x^2, right?

OpenStudy (anonymous):

then it becomes 2/x ln x^2?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

that is the derivative of the original expression...

OpenStudy (anonymous):

du/u is implicitly du/dx /u, since du is taken wrt x

OpenStudy (anonymous):

so i'm right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

alright, thanks!

OpenStudy (anonymous):

d/dx (ln(ln(x^2))=2/(x*ln(x^2))